IDVI Embedded Applet Test

This document was written in L^{A}TEX and processed for presentation as a web page by IDVI. The applet below was included by using a natural extension of the HyperTEX specials.

The applet tags placed in the original .tex file enclose alternate text, much as applet tags in an html file do. This text would be visible if the resulting dvi file were viewed with xdvi. If the width and height tags for the applet had been omitted, then the applet would have been sized to exactly cover this enclosed, invisible, text.

In the example above, the applet tags enclose a horizontal rule which is the width of the page and one pixel high. If both width and height had been omitted, we would have had a very short applet. Instead, the height was specified to be 150 pixels. The result is that, at every scale factor, the applet will be as wide as the rule which it encloses and 150 pixels high.